OpenStudy (anonymous):

how do i differenciate 2sin^3 (2x)?

OpenStudy (anonymous):

2*3 (COS^2(2X)*(2)

OpenStudy (anonymous):

\[\huge 12\cos(2x).(\sin(2x))^2\]

OpenStudy (anonymous):

apply chain rule

OpenStudy (anonymous):

if only i knew how to do that lol

OpenStudy (anonymous):

says that if g is a function that is differentiable at a point c (i.e. the derivative g′(c) exists) and f is a function that is differentiable at g(c), then the composite function f ∘ g is differentiable at c, and the derivative is

OpenStudy (anonymous):

\[(a*b)\prime=a \prime*b+b \prime*a\]

OpenStudy (anonymous):

(f∘ g)'(c)=f'(g(t))g'(t)